The mean salary offered to students who are graduating from Coastal State University this year is $24,245, with a standard deviation of $3712. A random sample of 85 Coastal State students graduating this year has been selected. What is the probability that the mean salary offer for these 85 students is $23,500 or more?
Carry your intermediate computations to at least four decimal places. Round your answer to at least three decimal places.

Answer by stanbon(75887)

The mean salary offered to students who are graduating from Coastal State University this year is $24,245, with a standard deviation of $3712. A random sample of 85 Coastal State students graduating this year has been selected. What is the probability that the mean salary offer for these 85 students is $23,500 or more?
------
z(23,500) = (23500-24245)/[3712/sqrt(85)] = -1.8504
-----
P(x-bar >= 23500) = P(z >= -1.8504) = 0.9679
=========================
Cheers,
Stan H.
--------------------------



Carry your intermediate computations to at least four decimal places. Round your answer to at least three decimal places.